The foot of a ladder is placed 12 feet from a wall. If the top of the ladder rests 9 feet up on the wall, how long is the ladder?
Answer by nerdybill(7384)

The foot of a ladder is placed 12 feet from a wall. If the top of the ladder rests 9 feet up on the wall, how long is the ladder?
.
Apply Pythagorean theorem:
Let x = length of ladder
then
9^2 + 12^2 = x^2
81 + 144 = x^2
225 = x^2
15 feet = x